Prev Next

SW&IT Meeting Minutes 2018-04-16

Attendees: Andrew, David, Doug, Igor, Joe, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang

Previous Action Items

  • NetBrain (David - schedule demo)
  • VM/SSD (David - followup meeting)
  • Scriptor follow-up (Igor)
  • Snapshot Argos tool (Matthieu) missing doc, howto, experts only ? Tomaso who is leaving ? Doug Miller

Outlook

News from management

  • board meeting Mon/Tue
  • monthly reports
  • time accounting

Team reports

AC

Last week
  • Added new commands to save the last LAT positions as nominal
  • Added mechanism to start/stop LAT spiral search loop
  • Added spiral search loop status to summary
This week
  • ARGOS run
    • test LAT spiral search
    • test sharpen algorithm for LGS acquisition
    • test find lasers algorithm for DX
    • review binocular arbitrator status GUI
    • review changes to arbitrator GUIs and others
    • test PCS coordinates when SlewToHold flag is used
    • review snapshot analyzer

DE

  • Veeam licensing, etc as predicate to moving VMs to local storage
    • Set up scheduled backups
    • Test restore
  • Getting Laura's laptop sorted
  • Netbrain demo; should be quick--what time?

DS

Last Week:
  • GCS / PCS acquisition handshake logging review (done + acquisitions log parsing tool development)
  • ARGOS pointing review (backtest & prep for run) - last run data reviewed, but not modeled or backtested)
This Week:
  • ARGOS pointing models & backtests in prep for upcoming run
  • Some support to Matthieu on mgt transition tasks

IS

Last week:
  • I spend most the week working mostly on PIT and some OT issues/releases.
  • For the PIT we still have/wait for the Italians partners to respond to some questions or even more request. One the other hand some or all of the work we\x92ve done might not go anywhere.
This week:
  • Will work on the getting the summary status for the LUCI scheduler as part of IT#6138

JA

Last week
  • Planner Config issue resolved.
  • Agree upon Subversion Queue revision control system charter (JA, ME, MB, IS).
  • IS requested that the OT Scripting UI have scrollbars so the contents are never crushed beyond readability. DONE for all instruments, svn revision r2747.
    • OT 2018A.1.3.0 released
  • Meetings with ME - Planner (demo, Synchronous Offsets)
This week
  • Planner - Implement Synchronous Offsets
  • Test OT Masks?

KS

Last week
  • Updated AO ALH config file for Xianyu's requested clarifications.
  • Integrated EPICS ezca into OAC HK upgrade - now it writes the temps/voltages from the HK board to EPICS channels for alarming.
    See Software/OACHousekeepingUpgrade#Software_Status for more detailed status.

This week
  • OAC HK upgrade:
    • handle errors from EPICS writes
    • add the harder alarming cases - if it's off, but wasn't commanded by OAC to be off
    • fix the argument options for the client programs
    • make a "poweronly" option for the getdata 8051 board

MB

  • management OC, WPC
  • techtalk#1 git//ci/rpm/kickstart
  • automate git/github support-tool, tcs/conf (IT x,y)
  • study INDI
  • 'queue' repo guidelines: stable trunk, feature branches
  • review `tools`
  • meeting Argos, Logging, rpm, telescope home, TCS configs, rcs/backup

Next

  • Argos, iLocater follow-up with partners
  • mountain trip, meet Leroy, Chris, discuss summit computers, fiber connectivity, Argos night
  • git migration kickoff
  • responsibilities and assignments draft

PK

  • completed and close dome tested IT 6746 (NS jumps)
  • fixed IT 7131 (ARGSO SAS and slewToHold)
  • upgrades to telemetry-graph tool
  • IT 7190 (LSS backlog, too many CLOSE_WAIT sockets), started lbto.logs Python library to parse TCS logs
  • closed IT 7176 (mcspu log rotation), IT 6978 (jet emails)
  • installed vim tagbar ( https://majutsushi.github.io/tagbar/) on my dev machine

This week:
  • 7190 will probably need redesign of RPC calls (reusing TCP/IP sockets)
  • PCS dies (various ITs)
  • tpk TimeToLimit resolution (IT 3949)

SH

XZ

Lat week:
  • Updated Alarm handler for ASM with Kellee
  • DX fastlink changed for LBTI
  • Meeting ARGOS, LBTI, AO
  • Summer adaptive secondary plan discussion with Guido and Mike L.
  • Sent power black plane test in Tucson requirements to MG
  • FLAO/UAO compile and build discussion with Yang
  • Changed the AOS and its GUI for the power backplane information
  • Changed the AOS and its GUI for the remove the jumping actuator
  • Validated new TT LUT for ARGOS
  • Testing the changed AOS with changed FLAO/UAO on both simulators
This week:
  • Fix the ice issue between AOS UAO by simulators
  • Continue working on the power backplane
  • FLAO/ ARGOS TT LUT support

YZ

  • Try to rebuild the "mergelogs" tool based on standard default versions of the the Boost library (and other dependencies). This will involve modifications to source code as well. This will provide a path for an automated building/deploying process.
  • Instead of building the AO code on a new VM, I will try to build the existing AO VM with newer versions of GCC (GCC 4.8+). The plan is to install Redhat SCL on the AO VM and use it to switch to different GCC versions. If succeeded, this will pave the way for the AO code to use C++11 features, which will greatly expand the options available.
  • Investigating ways to implement custom special sink in the "spdlog" library that sorts multiple async logging entries according to the log entry timestamp, and then forward the sorted log entries to other standard sinks.

Action items

-- %USERSIG{MatthieuBec}%

* svn2git:
---++ Comments Screen_Shot_2018-04-16_at_9.47.23_AM.png -- %BUBBLESIG{JosephAstier - 2018-04-11}%
 

Topic revision: r23 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback